Kilifi Ltd has 3.2 million shares and issue. The market price per share is sh. 75....

  1. Kilifi Ltd has 3.2 million shares and issue. The market price per share is sh. 75. The management of the company is contemplating raising an additional sh. 120 million through a rights issue in order to fund a new project. The right issue would be made at a discount of 20% of the current market price.

Required

  1. The number of shares to be issued in the rights issue                            (2 marks)
  2. The theoretical ex-right price- the value of a right                                            (5 marks)
Answer #1

I number of shares to issue = amount to raise/(current price*(1-discount %) = 120/(75*(1-0.2) = 2m

ii

Ex right price = (current share*current price+amount to raise)/(current shares+number of shares issued)

=(75*3.2+120)/(3.2+2) =

69.23